# Towards Minimizing Feature Drift in Model Merging: Layer-wise Task Vector Fusion for Adaptive Knowledge Integration

### v1

@@ C.3 Baseline Methods
  • EMR Merging (Huang et al., 2024): This method applies lightweight task-specific masks and rescalers to compress task vectors.
  • WEMOE (Tang et al., 2024): Upscales MLP layers into input-dependent Mixture-of-Experts modules to dynamically integrate shared and task-specific knowledge during model merging.
@@
  • AdaMerging (Yang et al., 2024b): This method adapts the model to new tasks by using an unlabeled test set to learn adaptive merging coefficients at both the task- and layer-level, as applied in Task Arithmetic.
  • AdaMerging++ (Yang et al., 2024b): An improved version of AdaMerging, which integrates the principles of Ties-Merging (Yadav et al., 2023) to further enhance the adaptive merging process.
  • Surgery (Yang et al., 2024a): Surgery introduces a feature transformation module that aligns the features from different tasks during the merging process. In our experiments, we utilize the basic version of Surgery alongside Task Arithmetic for evaluation.
  • Localize-and-Stitch (He et al., 2025): This method optimally combines the strengths of several fine-tuned models by identifying and localizing essential regions within each model before merging.
@@
  • Fisher Merging (Matena & Raffel, 2022): This approach uses the Fisher information matrix to assess the relative importance of model parameters, merging them based on their significance.
  • RegMean (Jin et al., 2023): RegMean adjusts and combines rows from weight matrices based on statistical information gathered from the training data, refining the model’s parameters.
  • Task Arithmetic (Ilharco et al., 2023): Task Arithmetic introduces the concept of a “task vector,” which is defined as the difference between fine-tuned model parameters and pre-trained model parameters. By combining multiple task vectors and adding them to the pre-trained model, it enables multi-task learning.
  • Ties-Merging (Yadav et al., 2023): Ties-Merging removes insignificant parameters from the task vectors and resolves sign conflicts, reducing interference during the merging of task vectors.
  • TATR (Sun et al., 2025b): TATR improves upon Task Arithmetic by restricting the merging of task vectors to a defined trust region, which reduces knowledge conflicts between tasks.
  • TATR & Ties-Merging (Sun et al., 2025b; Yadav et al., 2023): This method combines the trust region restriction of TATR with Ties-Merging to further enhance task vector merging.
  • Consensus Merging (Wang et al., 2024): This method computes a set of masks for each task vector to minimize the distance between the merged model and each fine-tuned model in the parameter space.
  • AWD Merging (Xiong et al., 2024): AWD Merging generates redundant vectors such that subtracting them from the original task vectors leads to increased orthogonality in the remaining vectors.
  • PCB Merging (DU et al., 2024): PCB Merging trims components of the task vector that have small magnitudes and are not strongly correlated with other tasks.
  • CAT Merging (Sun et al., 2025a): CAT Merging trims components (using projection or mask) of the task vector that may cause knowledge conflicts.
@@ D.1 Additional Performance Comparision
In this section, we compare LOT Merging with both compression-based methods and test-time adaptation-based approaches. Compression-based methods rely on techniques such as masking (Huang et al., 2024) or singular value decomposition (SVD) (Tang et al., 2024) to compress task vectors, and typically require manually specified or trained routers during inference to select the appropriate task vector. On the other hand, test-time adaptation-based methods leverage unlabeled test data to train merging weights or certain modules, which introduces additional computational overhead and data requirements.
@@ D.6 Analysis of Computational Complexity
  • Optimal task‐vector computation: This step relies on a closed-form solution, eliminating the need for iterative optimization. For linear task representations, this involves a matrix inversion operation. While matrix inversion can be associated with higher theoretical complexity ($\mathcal{O}(d^{3})$), we mitigate this overhead through a parallelized implementation on modern GPUs. In practice, the computation is highly efficient and introduces negligible latency.
Empirical runtime measurements, summarized in Table 9 (recorded on a single NVIDIA RTX 3090 GPU), confirm that LOT Merging achieves significantly lower wall-clock time compared to training-based baselines such as TA with Surgery (Yang et al., 2024a) and AdaMerging (Yang et al., 2024b). These results highlight the practical efficiency of our approach, making it highly suitable for scenarios where rapid adaptation across tasks is required.

### v2

@@ C.3 Baseline Methods
  • EMR Merging (emrmerging, ): This method applies lightweight task-specific masks and rescalers to compress task vectors.
  • WEMOE (wemoe, ): Upscales MLP layers into input-dependent Mixture-of-Experts modules to dynamically integrate shared and task-specific knowledge during model merging.
@@
  • AdaMerging (adamerging, ): This method adapts the model to new tasks by using an unlabeled test set to learn adaptive merging coefficients at both the task- and layer-level, as applied in Task Arithmetic.
  • AdaMerging++ (adamerging, ): An improved version of AdaMerging, which integrates the principles of Ties-Merging (tiesmerging, ) to further enhance the adaptive merging process.
  • Surgery (surgerymerging, ): Surgery introduces a feature transformation module that aligns the features from different tasks during the merging process. In our experiments, we utilize the basic version of Surgery alongside Task Arithmetic for evaluation.
  • Localize-and-Stitch (localizeandstitch, ): This method optimally combines the strengths of several fine-tuned models by identifying and localizing essential regions within each model before merging.
@@
  • Fisher Merging (fisher_merging, ): This approach uses the Fisher information matrix to assess the relative importance of model parameters, merging them based on their significance.
  • RegMean (regmean, ): RegMean adjusts and combines rows from weight matrices based on statistical information gathered from the training data, refining the model’s parameters.
  • Task Arithmetic (task_arithmetic, ): Task Arithmetic introduces the concept of a “task vector,” which is defined as the difference between fine-tuned model parameters and pre-trained model parameters. By combining multiple task vectors and adding them to the pre-trained model, it enables multi-task learning.
  • Ties-Merging (tiesmerging, ): Ties-Merging removes insignificant parameters from the task vectors and resolves sign conflicts, reducing interference during the merging of task vectors.
  • TATR (tatr, ): TATR improves upon Task Arithmetic by restricting the merging of task vectors to a defined trust region, which reduces knowledge conflicts between tasks.
  • TATR & Ties-Merging (tatr, ; tiesmerging, ): This method combines the trust region restriction of TATR with Ties-Merging to further enhance task vector merging.
  • Consensus Merging (consensus_merging, ): This method computes a set of masks for each task vector to minimize the distance between the merged model and each fine-tuned model in the parameter space.
  • AWD Merging (awdmerging, ): AWD Merging generates redundant vectors such that subtracting them from the original task vectors leads to increased orthogonality in the remaining vectors.
  • PCB Merging (pcbmerging, ): PCB Merging trims components of the task vector that have small magnitudes and are not strongly correlated with other tasks.
  • CAT Merging (catmerging, ): CAT Merging trims components (using projection or mask) of the task vector that may cause knowledge conflicts.
@@ D.1 Additional Performance Comparision
In this section, we compare LOT Merging with both compression-based methods and test-time adaptation-based approaches. Compression-based methods rely on techniques such as masking emrmerging or singular value decomposition (SVD) wemoe to compress task vectors, and typically require manually specified or trained routers during inference to select the appropriate task vector. On the other hand, test-time adaptation-based methods leverage unlabeled test data to train merging weights or certain modules, which introduces additional computational overhead and data requirements.
@@ D.6 Analysis of Computational Complexity
  • Optimal task‐vector computation: This step relies on a closed-form solution, eliminating the need for iterative optimization. For linear task representations, this involves a matrix inversion operation. While matrix inversion can be associated with higher theoretical complexity ($\mathcal{O}(d^{3})$), we mitigate this overhead through parallelized implementation on modern GPUs. In practice, the computation is highly efficient and introduces negligible latency.
Empirical runtime measurements, summarized in Table 9 (recorded on a single NVIDIA RTX 3090 GPU), confirm that LOT Merging achieves significantly lower wall-clock time compared to training-based baselines such as TA with Surgery surgerymerging and AdaMerging adamerging . These results highlight the practical efficiency of our approach, making it highly suitable for scenarios where rapid adaptation across tasks is required.
@@ D.7 Analysis of Consistency
Table 11: Results (%) on seen and unseen Tasks
SUN397 Cars RESISC45 DTD SVHN GTSRB Avg Acc (Seen) MNIST (Unseen) EuroSAT (Unseen) Avg Acc (Unseen)
Task Arithmetic 63.3 62.4 75.1 57.8 84.6 80.4 70.6 77.2 46.2 61.7
Ties-Merging 67.8 66.2 77.2 56.7 77.1 70.9 69.3 75.9 43.3 59.6
LOT Merging 69.1 69.1 89.7 65.6 94.1 93.7 80.2 81.5 45.4 63.5
SUN397 Cars GTSRB EuroSAT DTD MNIST Avg Acc (Seen) RESISC45 (Unseen) SVHN (Unseen) Avg Acc (Unseen)
Task Arithmetic 64.0 64.0 75.2 87.7 57.0 95.7 73.9 52.3 44.9 51.1
Ties-Merging 68.0 67.1 67.7 78.4 56.5 92.8 71.8 58.7 49.2 53.9
LOT Merging 68.8 69.9 84.8 95.8 61.5 97.5 79.7 56.3 54.6 55.5
Table 12: Average accuracy with Gaussian blur kernel
Method Gaussian Blur Kernel ViT-B/32 (%) ViT-L/14 (%)
Task Arithmetic 69.1 84.5
LOT Merging $kernel\_size=3,\sigma=1$ 82.2 90.3
LOT Merging $kernel\_size=5,\sigma=2$ 80.0 89.5
LOT Merging $kernel\_size=7,\sigma=3$ 78.3 88.7
Table 13: Performance when using similar distributed exemplars
Method COCO Caption OKVQA ScienceQA
Task Arithmetic 0.53 40.37 51.55
LOT Merging 0.87 44.25 60.08
LOT Merging w/ surrogate exemplars 0.62 40.15 53.54
D.8 Analysis of Generalization Ability
We first evaluated our method by merging models trained on six out of eight tasks and then testing on the remaining two unseen tasks. As shown in the tables 11, LOT Merging consistently outperforms both Task Arithmetic and Ties-Merging—not only on seen tasks, but also on unseen (out-of-domain) tasks. These results demonstrate the moderate generalization ability of our approach. However, the performance gap does become narrower on out-of-domain tasks compared to in-domain tasks. This observation motivates us to further explore ways to enhance LOT Merging’s effectiveness on out-of-domain tasks in future work.
Next, we explored the use of domain-similar auxiliary data by applying a Gaussian blur to the original task exemplars. As shown in Table 12, LOT Merging consistently outperformed Task Arithmetic across all blur levels, although performance gradually declined as blur strength increased.
To investigate whether similar distribution exemplars would benefit, we merged BLIP models fine-tuned on COCO Caption, OKVQA, and ScienceQA, but replaced each task’s exemplars with samples from semantically related datasets: Flickr30k for COCO Caption, AOKVQA for OKVQA, and IconQA for ScienceQA. As shown in Table 13, LOT Merging with surrogate exemplars outperforms Task Arithmetic, though the improvements are task-dependent—substantial for COCO Caption, but more limited for the other two.
D.9 Robustness Analysis with Data Corruption
To assess the robustness of LOT Merging under noisy conditions, we adopted the corruption protocol introduced in adamerging , applying seven types of corruption to both exemplars and test sets. We evaluated the performance by merging eight ViT-B/32 models on each corrupted dataset and compared LOT Merging to Task Arithmetic. Table 14 shows that LOT Merging maintains consistent improvement over Task Arithmetic under all corruption types.
Table 14: Average accuracy under different corruption types
Corruption Type Task Arithmetic LOT Merging (↑ $\Delta$)
Clean (no corruption) 69.1 82.7 (↑ 13.6)
Motion Blur 47.3 59.9 (↑ 12.6)
Impulse Noise 51.1 60.8 (↑ 9.7)
Gaussian Noise 50.6 60.1 (↑ 9.5)
Pixelate 48.9 61.8 (↑ 12.9)
Spatter 65.0 79.2 (↑ 14.2)
Contrast 38.5 45.9 (↑ 7.4)
JPEG Compression 25.9 28.1 (↑ 2.2)
Table 15: Performance of Different Algorithms on GLUE Tasks
Algorithm CoLA MNLI MRPC QNLI QQP RTE SST-2 STS-B Average #best
Task Arithmetic 6.68 66.23 78.46 78.62 72.69 53.43 83.49 27.10 58.34 1
Ties-Merging 9.46 59.34 74.71 65.93 41.29 47.29 72.13 9.21 47.42 0
PCB Merging 11.40 50.85 77.63 78.22 55.78 60.29 75.57 67.01 59.59 1
CAT Merging 33.20 72.33 68.22 82.92 76.05 62.82 89.33 15.57 62.56 2
LOT Merging 17.13 73.04 78.27 77.22 78.81 65.35 89.74 25.50 63.13 4
Table 16: Performance Comparison when merging two Qwen models.
Method ArguAna ArXiv Hierarchical Clustering S2S
Individual 53.62 0.6377
Task Arithmetic 30.86 0.6318
LOT Merging 47.36 0.6374
D.10 Results on Merging Language Models
We conducted experiments using RoBERTa roberta as the backbone on the GLUE benchmark glue , which comprises eight diverse NLP tasks spanning both classification and regression (e.g., stsb). For evaluation, we report accuracy for classification tasks and the average of Pearson and Spearman correlations for the regression task. The results are presented in Table 15. LOT Merging achieves the highest average performance compared to the baseline methods, demonstrating its effectiveness for language model merging.
We also conduct experiments on large language models. We select Qwen3-0.6B qwen3 as the pretrained backbone, and two official fine-tuned variants—Qwen3-Embedding-0.6B and Qwen3-Reranker-0.6B. Although Qwen3-0.6B has a relatively small parameter count, it features a deep architecture with 28 Qwen3DecoderBlocks, each comprising an attention module (four projection linear layers) and an MLP module (three projection linear layers), for a total of 196 linear layers.
To reflect their respective strengths in embedding and ranking, we selected two tasks from the MTEB (English, v2) benchmark mteb : ArguAna — a retrieval task requiring strong ranking capabilities; ArXiv Hierarchical Clustering S2S — a clustering task requiring high-quality embeddings. Evaluation metrics were top-1 accuracy for ArguAna and V-measure for ArXiv Hierarchical Clustering. Results in Table 16 demonstrate that LOT Merging substantially outperforms Task Arithmetic on ArguAna, indicating improved cross-task transfer for ranking, while maintaining competitive clustering performance. These results confirm that our method remains effective on deep LLMs such as Qwen3.
Table 17: Pearson Correlation across Different Tasks
COCO Caption Flickr30k Caption TextCaps OKVQA TextVQA ScienceQA
Pearson Correlation 0.51 0.60 0.30 0.82 0.77 0.20
D.11 Analysis of Feature Drift on Multi-Model Tasks
To assess the impact of feature drift in vision-language settings, we measured the Pearson correlation between feature drift and loss change in BLIP models across several multimodal benchmarks after Task Arithmetic merging. The results in  17 demonstrate that feature drift remains a positively correlated factor.
While the correlation strength varies—reflecting inherent complexities such as reasoning demands and modality interactions—our experiments consistently show that reducing feature drift leads to performance gains, even in vision-language tasks. This suggests that, despite the greater complexity in multimodal models, minimizing feature drift remains an effective and broadly applicable merging objective.
Table 18: Effect of Exemplar Classes on Average Accuracy
Backbone Method Exemplar Classes Avg. Acc.
ViT-B/32 Task Arithmetic 69.1
ViT-B/32 LOT Merging All classes 82.7
ViT-B/32 LOT Merging First 1/2 81.8
ViT-B/32 LOT Merging First 1/3 81.0
ViT-B/32 LOT Merging First 1/4 79.6
ViT-B/32 LOT Merging First 1/5 77.6
ViT-L/14 Task Arithmetic 84.5
ViT-L/14 LOT Merging All classes 90.5
ViT-L/14 LOT Merging First 1/2 90.0
ViT-L/14 LOT Merging First 1/3 89.6
ViT-L/14 LOT Merging First 1/4 88.9
ViT-L/14 LOT Merging First 1/5 88.3
D.12 Robustness Analysis under Real-Time Data Distribution
We further study the impact of real-time data on model performance, particularly in scenarios where exemplars are available for only a subset of classes, while evaluation is conducted on the complete test set. The results in Table 18 show that although LOT Merging’s performance declines with fewer exemplar classes, it still consistently outperforms Task Arithmetic.